An investor is considering purchasing a 20-year 7% coupon bond selling for $816 and a par value of $1,000. Calculate the interest on interest from the bond assuming that the semi-annual coupon payments can be reinvested at 4½% every six months.
Answer #1
Semiannual coupons = 1000*3.5% = $ 35.00
FV of the reinvested coupons = 35*(1.045^40-1)/0.045 = $       3,746.06
Interest on interest = 3746.06-35*40 = $       2,346.06